# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ 2.2 The power of the $\ell$-test
To explain this behavior, we first characterize the relationship between the $t$-test statistic and $u_{1}$ in Lemma 2.2. It turns out that $u_{1}$ is a scalar multiple of $\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{y}$ (we argue this in Equations (C.1) through (C.4) of the Appendix and is a major component in the proof of the lemma), and hence, $u_{1}$ is a measure of the association between $\bm{X}_{j}$ and the component of $\bm{y}$ that cannot be explained by the rest of the columns.
Lemma 2.2.
@@
We prove this result in Section C.2 of the appendix. Lemma 2.2 and Theorem 2.1 together show that a (one-sided) conditional-on-$\bm{S}^{(j)}$ test based on any of the test-statistics—$u_{1},T_{j}$ and $\hat{\beta}_{j}^{\lambda}$, yield exactly the same p-values as long as the observed LASSO estimate is non-zero, as in this case all the three test statistics are strictly increasing in each other. We also know that $T_{j}$ is independent of $\bm{S}^{(j)}$, which follows from standard theory on ancillary statistics, however we supply a separate proof for this in Section C.4 of the Appendix. This implies that a one-sided conditional test based on $T_{j}\mid\bm{S}^{(j)}$ yields exactly the one-sided $t$-test p-value, which by the above argument is exactly equal to the one-sided p-value of the conditional test based on $\hat{\beta}^{\lambda}_{j}\mid\bm{S}^{(j)}$ when the observed LASSO estimate is non-zero.
@@
Even though the above paragraph establishes that conditional one-sided testing based on $\hat{\beta}^{\lambda}_{j}$ can do only as well as the corresponding one-sided $t$-test, we will now argue that the former test can gain considerable power over the $t$-test in the two-sided testing regime. As a first step, note that we can use Theorem 2.1 to characterize the set :=$\mathcal{R}:=\{u_{1}:\hat{\beta}_{j}^{\lambda}\neq 0\}$ as a disjoint union of two intervals: $\mathcal{R}=(-\infty,v_{-})\cup(v_{+},\infty)$, where, $v_{\pm}=\Lambda_{j}(0,\pm 1)$.
@@
Figure 1: Connection between the distribution of $u_{1}$ (left) and the $\ell$-distribution (right). Regions of matching color between the two plots represent a one-to-one correspondence between the random variable values within them, as proved in Theorem 2.1.
@@
We will argue that when $\beta_{j}\neq 0$, the test based on $|\hat{\beta}^{\lambda}_{j}|$ (i.e., the $\ell$-test) leverages the asymmetry of the interval $[v_{-},v_{+}]$ about 0 to gain power. Observe that from Theorem 2.1, $\hat{\beta}_{j}^{\lambda}$ is negative when $u_{1}\leq v_{-}$, while it is positive when $u_{1}\geq v_{+}$. Without loss of generality, we will assume that $\beta_{j}>0$ for the proceeding discussion, and assume the center of the interval $[v_{-},v_{+}]$ is negative (we will justify this latter assumption in a little bit). Consider Figure 1 for a visual representation of this, where under $H_{j}$, the left and the right figures show the conditional distributions of $u_{1}$ (i.e., $F_{u}$) and $\hat{\beta}_{j}^{\lambda}$ (i.e., the $\ell$-distribution $F_{\ell}^{\lambda}(\cdot\mid\bm{S}^{(j)})$), respectively. The correspondence between the two distributions is shown by matching colors—for example, as is evident from Theorem 2.1, the mass that the distribution of $u_{1}$ puts to the left of $v_{-}$ is exactly the mass the $\ell$-distribution puts on the negative half, and hence both these regions are colored green. As can be seen from the figure, the asymmetry in $[v_{-},v_{+}]$ (and the symmetry of $F_{u}$) directly implies asymmetry in the $\ell$-distribution.
@@
Since $\beta_{j}>0$, we expect that $\hat{\beta}_{j}^{\lambda}>0$ as well, as reflected in the righthand plot, with corresponding positive $u_{1}$ also marked in the lefthand plot. Due to the symmetry of $u_{1}$’s distribution, the p-value of the two-sided test using $|u_{1}|$ is just twice the mass to the right of $u_{1}$, and by Lemma 2.2, this is also the p-value of the (two-sided) $t$-test. To understand the $\ell$-test p-value for comparison, note that by Theorem 2.1, the mass to the right of $u_{1}$ in the left plot is exactly the mass to the right of $\hat{\beta}^{\lambda}_{j}$ in the right plot, but, critically, the mass to the left of $-\hat{\beta}^{\lambda}_{j}$ in the right plot is far less than the mass to the left of $-u_{1}$ in the left plot. Thus, the $\ell$-test’s p-value, which is exactly the mass of the shaded regions in the right plot, is dominated by the right-most shaded region, whose mass is exactly the value of the one-sided $t$-test.
@@
Next, we argue why we expect the interval $[v_{-},v_{+}]$ to lean opposite to the true sign of $\beta_{j}$, that is, towards the negative side in this case. Note that the interval $[v_{-},v_{+}]$ has mid-point
@@
Thus under the alternative, $m_{j}$’s distribution is shifted towards the opposite sign of $\beta_{j}$ as long as $\bm{X}_{j}$ is not exactly orthogonal to $\bm{X}_{-j}$. And when $\bm{\beta}$ is sparse, we expect the lasso estimator $\hat{\bm{\beta}}_{-j}^{\lambda}(0)$ of $\bm{\beta}_{-j}$ to be a good one, and hence that $\hat{m}_{j}$’s distribution will also be shifted towards the opposite sign of $\beta_{j}$. In particular, when $\beta_{j}>0$, this means we expect $[v_{-},v_{+}]$ to be shifted in the negative direction, as we assumed it would be earlier in this subsection.
In sum, when $\bm{\beta}$ is sparse, the LASSO leverages information in $\bm{S}^{(j)}$ (via $\hat{\bm{\beta}}_{-j}^{\lambda}(0)$) to guess the sign of $\beta_{j}$, and the point mass in the $\ell$-distribution uses that guess (via $[v_{-},v_{+}]$) to reduce the “wrong” tail of the $\ell$-test, resulting in a test with power approximating the one-sided $t$-test; see Appendix B for further discussion. Note that although the intuition for the power gain of the $\ell$-test over the $t$-test relied on sparsity, we emphasize that the validity guarantees of the $\ell$-test remain identical to those of the $t$-test, and in particular do not require sparsity.
@@ 2.3 Breaking ties when $\hat{\beta}_{j}^{\lambda}=0$
As alluded to earlier in this section, the $\ell$-test p-value $\bar{F}_{|\ell|}^{\lambda}(|\hat{\beta}^{\lambda}_{j}|\mid\bm{S}^{(j)})$ is not $\mathrm{Unif}(0,1)$ under $H_{j}$. Instead, its conditional distribution given $\bm{S}^{(j)}$ is a mixture of $\mathrm{Unif}(0,\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\neq 0\mid\bm{S}^{(j)}))$ and a point mass at 1 of weight $\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}=0\mid\bm{S}^{(j)})$ because $|\hat{\beta}_{j}^{\lambda}|=0$ is the “least significant” value of the test statistic and occurs with positive probability. It is preferable not to have such a point mass, since it makes the $\ell$-test somewhat conservative and because both users and many procedures which take p-values as inputs generally assume uniform null p-values. To remedy this, we need a way to break ties among data values that give $\hat{\beta}_{j}^{\lambda}=0$, since the test statistic $|\hat{\beta}_{j}^{\lambda}|$ does not distinguish between them. The strong connection between $\hat{\beta}_{j}^{\lambda}$ and $u_{1}$ established in Theorem 2.1 and visualized in Figure 1 suggests that $u_{1}$, whose distribution is continuous on $[v_{-},v_{+}]$ given the event $\{\hat{\beta}_{j}^{\lambda}=0\}$, provides a way forward. In particular, since $\hat{\beta}_{j}^{\lambda}$ approaches zero as $u_{1}$ approaches $v_{-}$ from the left or $v_{+}$ from the right, it is natural (and continuous in the data) to set $v_{-}$ and $v_{+}$ as tied for the “most significant” values on the interval $[v_{-},v_{+}]$, and then have the significance decrease as $u_{1}$ moves inward from those endpoints. This corresponds to breaking ties according to $|u_{1}-\hat{m}_{j}|$ when $\hat{\beta}_{j}^{\lambda}=0$, and can equivalently be thought of as using $|\hat{\beta}_{j}^{\lambda}|+\min\{|u_{1}-\hat{m}_{j}|-(v_{+}-\hat{m}_{j}),0\}$ as the test statistic. Recalling that $F_{u}$ denotes $u_{1}$’s CDF and defining $\bar{F}_{|u|}(u^{\prime}\mid\bm{S}^{(j)}):=\mathbb{P}_{H_{j}}(|u_{1}-\hat{m}_{j}|\geq u^{\prime}\mid\bm{S}^{(j)})=1-F_{u}(\hat{m}_{j}+u^{\prime})+F_{u}(\hat{m}_{j}-u^{\prime})$ as the tail probability of $|u_{1}-\hat{m}_{j}|$, we can express this p-value as
@@ 2.4 The choice of $\lambda$
Thus far, we have treated $\lambda$ as a fixed tuning parameter, but in practice it is preferable to have an automated, data-dependent way to choose it. Standard practice for the LASSO is to choose $\lambda$ via cross-validation (on the full data $(\bm{y},\bm{X})$), but while this choice invalidates the theoretical guarantees of the $\ell$-test, just a slight modification of it is sufficient to retain those guarantees. Let $\tilde{\bm{u}}\sim\mathrm{Unif}(\mathbb{S}^{n-d})$ be drawn independently of $\bm{u}$, and plug it into Equation (2.1) and call the resulting lefthand side $\tilde{\bm{y}}$, so that $\tilde{\bm{y}}$ is conditionally independent of $\bm{y}$ given $\bm{S}^{(j)}$. Then it is easy to see that cross-validation on $(\tilde{\bm{y}},\bm{X}_{-j})$ produces a $\lambda$, which we denote by $\hat{\lambda}_{j}$, that is exactly valid to use in the $\ell$-test, since conditioning on $\hat{\lambda}_{j}$ does not change the $\ell$-distribution. [^2] Cross-validation on $(\tilde{\bm{y}},\bm{X})$ would also be valid and natural, but we prefer $(\tilde{\bm{y}},\bm{X}_{-j})$ for computational reasons; see Appendices E.1 and E.2 for details on this and other ways to choose $\lambda$ that we considered, all of which we found to be empirically dominated by $\hat{\lambda}_{j}$. And empirically, $\hat{\lambda}_{j}$ (our implementation uses 10-fold cross-validation) seems to be just as powerful as the (technically invalid) choice via cross-validation on $(\bm{y},\bm{X})$; see Appendix E.2. Although $\hat{\lambda}_{j}$ is randomized through $\tilde{\bm{u}}$ and the random 10-fold partition of the data used by cross-validation, we find this exogenous randomness barely makes a difference: in our simulations it empirically accounts for at most about 0.8% of the variability in the $\ell$-test p-value $p_{j}^{\hat{\lambda}_{j}}$; see Appendix E.2. Furthermore, if desired, this randomness could be arbitrarily reduced by computing many conditionally independent $\hat{\lambda}_{j}$’s and using their mean or median for the $\ell$-test.
@@ 2.5 Putting it all together: the $\ell$-test
We can now state our recommended implementation of the $\ell$-test: the p-value $p_{j}^{\hat{\lambda}_{j}}$ which combines the main $\ell$-test idea with the tie-breaking of Sections 2.3 and the $\lambda$ choice of Section 2.4. Computing $p_{j}^{\hat{\lambda}_{j}}$ requires, aside from a cross-validated LASSO to compute $\hat{\lambda}_{j}$, one (non-cross-validated) LASSO to compute $\hat{\bm{\beta}}^{\hat{\lambda}_{j}}$ and, if $\hat{\beta}_{j}^{\hat{\lambda}_{j}}\neq 0$, then a second (non-cross-validated) LASSO to compute $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(-\hat{\beta}_{j}^{\hat{\lambda}_{j}})$. When $\hat{\beta}_{j}^{\hat{\lambda}_{j}}\neq 0$, these two LASSO’s allow us to compute the two tails for the $\ell$-test p-value via Theorem 2.1, since $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(\hat{\beta}_{j}^{\hat{\lambda}_{j}})=\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}$. And when $\hat{\beta}_{j}^{\hat{\lambda}_{j}}=0$, by Equation (2.7), the only LASSO quantity needed is $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(0)$ to compute $\hat{m}_{j}$, but since $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(0)=\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}$ in this case, no additional LASSO run is needed beyond the first. Thus computation for the $\ell$-test requires just a very small constant number of LASSO runs.
@@
Corollary 2.1 (Validity of the $\ell$-test).

### v2

@@ 2.2 The power of the $\ell$-test
To explain this behavior, we first characterize the relationship between the $t$-test statistic and $u_{1}$ in Lemma 2. It turns out that $u_{1}$ is a scalar multiple of $\bm{X}_{j}^{T}(\bm{I}-\bm{P}_{-j})\bm{y}$ (we argue this in Equations (D.1) through (D.4) of the Appendix and is a major component in the proof of the lemma), and hence, $u_{1}$ is a measure of the association between $\bm{X}_{j}$ and the component of $\bm{y}$ that cannot be explained by the rest of the columns.
Lemma 2.
@@
[…]
@@
Even though the above paragraph establishes that conditional one-sided testing based on $\hat{\beta}^{\lambda}_{j}$ can do only as well as the corresponding one-sided $t$-test, we will now argue that the former test can gain considerable power over the $t$-test in the two-sided testing regime. As a first step, note that we can use Theorem 1 to characterize the set $\mathcal{R}:=\{u_{1}:\hat{\beta}_{j}^{\lambda}\neq 0\}$ as a disjoint union of two intervals: $\mathcal{R}=(-\infty,v_{-})\cup(v_{+},\infty)$, where, $v_{\pm}=\Lambda_{j}(0,\pm 1)$.
@@
Figure 1: Connection between the distribution of $u_{1}$ (left) and the $\ell$-distribution (right). Regions of matching color between the two plots represent a one-to-one correspondence between the random variable values within them, as proved in Theorem 1. This figure is for demonstration and does not necessarily reproduce the exact shapes of the $u_{1}$- or $\ell$-distributions.
@@
We will argue that when $\beta_{j}\neq 0$, the test based on $|\hat{\beta}^{\lambda}_{j}|$ (i.e., the $\ell$-test) leverages the asymmetry of the interval $v_{-},v_{+}]$ about 0 to gain power. Observe that from Theorem 1, $\hat{\beta}_{j}^{\lambda}$ is negative when $u_{1}\leq v_{-}$, while it is positive when $u_{1}\geq v_{+}$. Without loss of generality, we will assume that $\beta_{j}>0$ for the proceeding discussion, and assume the center of the interval $v_{-},v_{+}]$ is negative (we will justify this latter assumption in a little bit). Consider Figure 1 for a visual representation of this, where under $H_{j}$, the left and the right figures show the conditional distributions of $u_{1}$ (i.e., $F_{u}$) and $\hat{\beta}_{j}^{\lambda}$ (i.e., the $\ell$-distribution $F_{\ell}^{\lambda}(\cdot\mid\bm{S}^{(j)})$), respectively. The correspondence between the two distributions is shown by matching colors—for example, as is evident from Theorem 1, the mass that the distribution of $u_{1}$ puts to the left of $v_{-}$ is exactly the mass the $\ell$-distribution puts on the negative half, and hence both these regions are colored green. As can be seen from the figure, the asymmetry in $v_{-},v_{+}]$ (and the symmetry of $F_{u}$) directly implies asymmetry in the $\ell$-distribution.
@@
Since $\beta_{j}>0$, we expect that $\hat{\beta}_{j}^{\lambda}>0$ as well, as reflected in the right-hand plot, with corresponding positive $u_{1}$ also marked in the left-hand plot. Due to the symmetry of $u_{1}$’s distribution, the $p$-value of the two-sided test using $|u_{1}|$ is just twice the mass to the right of $u_{1}$, and by Lemma 2, this is also the $p$-value of the (two-sided) $t$-test. To understand the $\ell$-test $p$-value for comparison, note that by Theorem 1, the mass to the right of $u_{1}$ in the left plot is exactly the mass to the right of $\hat{\beta}^{\lambda}_{j}$ in the right plot, but, critically, the mass to the left of $\hat{\beta}^{\lambda}_{j}$ in the right plot is far less than the mass to the left of $u_{1}$ in the left plot. Thus, the $\ell$-test’s $p$-value, which is exactly the mass of the shaded regions in the right plot, is dominated by the right-most shaded region, whose mass is exactly the value of the one-sided $t$-test.
@@
Next, we argue why we expect the interval $v_{-},v_{+}]$ to lean opposite to the true sign of $\beta_{j}$, that is, towards the negative side in this case. Note that the interval $v_{-},v_{+}]$ has mid-point
@@
Thus under the alternative, $m_{j}$’s distribution is shifted towards the opposite sign of $\beta_{j}$ as long as $\bm{X}_{j}$ is not exactly orthogonal to $\bm{X}_{-j}$. And when $\bm{\beta}$ is sparse, we expect the lasso estimator $\hat{\bm{\beta}}_{-j}^{\lambda}(0)$ of $\bm{\beta}_{-j}$ to be a good one, and hence that $\hat{m}_{j}$’s distribution will also be shifted towards the opposite sign of $\beta_{j}$. In particular, when $\beta_{j}>0$, this means we expect $v_{-},v_{+}]$ to be shifted in the negative direction, as we assumed it would be earlier in this subsection.
In sum, when $\bm{\beta}$ is sparse, the LASSO leverages information in $\bm{S}^{(j)}$ (via $\hat{\bm{\beta}}_{-j}^{\lambda}(0)$) to guess the sign of $\beta_{j}$, and the point mass in the $\ell$-distribution uses that guess (via $v_{-},v_{+}]$) to reduce the “wrong” tail of the $\ell$-test, resulting in a test with power approximating the one-sided $t$-test; see Appendix B for further discussion. Note that although the intuition for the power gain of the $\ell$-test over the $t$-test relied on sparsity, we emphasize that the validity guarantees of the $\ell$-test remain identical to those of the $t$-test, and in particular do not require sparsity.
@@ 2.3 Breaking ties when $\hat{\beta}_{j}^{\lambda}=0$
As alluded to earlier in this section, the $\ell$-test $p$-value $\bar{F}_{|\ell|}^{\lambda}(|\hat{\beta}^{\lambda}_{j}|\mid\bm{S}^{(j)})$ is not $\mathrm{Unif}(0,1)$ under $H_{j}$. Instead, its conditional distribution given $\bm{S}^{(j)}$ is a mixture of $\mathrm{Unif}(0,\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}\neq 0\mid\bm{S}^{(j)}))$ and a point mass at 1 of weight $\mathbb{P}_{H_{j}}(\hat{\beta}_{j}^{\lambda}=0\mid\bm{S}^{(j)})$ because $|\hat{\beta}_{j}^{\lambda}|=0$ is the “least significant” value of the test statistic and occurs with positive probability. It is preferable not to have such a point mass, since it makes the $\ell$-test somewhat conservative and because both users and many procedures which take $p$-values as inputs generally assume uniform null $p$-values. To remedy this, we need a way to break ties among data values that give $\hat{\beta}_{j}^{\lambda}=0$, since the test statistic $|\hat{\beta}_{j}^{\lambda}|$ does not distinguish between them. The strong connection between $\hat{\beta}_{j}^{\lambda}$ and $u_{1}$ established in Theorem 1 and visualized in Figure 1 suggests that $u_{1}$, whose distribution is continuous on $v_{-},v_{+}]$ given the event $\{\hat{\beta}_{j}^{\lambda}=0\}$, provides a way forward. In particular, since $\hat{\beta}_{j}^{\lambda}$ approaches zero as $u_{1}$ approaches $v_{-}$ from the left or $v_{+}$ from the right, it is natural (and continuous in the data) to set $v_{-}$ and $v_{+}$ as tied for the “most significant” values on the interval $v_{-},v_{+}]$, and then have the significance decrease as $u_{1}$ moves inward from those endpoints. This corresponds to breaking ties according to $|u_{1}-\hat{m}_{j}|$ when $\hat{\beta}_{j}^{\lambda}=0$, and can equivalently be thought of as using $|\hat{\beta}_{j}^{\lambda}|+\min\{|u_{1}-\hat{m}_{j}|-(v_{+}-\hat{m}_{j}),0\}$ as the test statistic. Recalling that $F_{u}$ denotes $u_{1}$’s CDF and defining $\bar{F}_{|u|}(u^{\prime}\mid\bm{S}^{(j)}):=\mathbb{P}_{H_{j}}(|u_{1}-\hat{m}_{j}|\geq u^{\prime}\mid\bm{S}^{(j)})=1-F_{u}(\hat{m}_{j}+u^{\prime})+F_{u}(\hat{m}_{j}-u^{\prime})$ as the tail probability of $|u_{1}-\hat{m}_{j}|$, we can express this $p$-value as
@@ 2.4 The choice of $\lambda$
Thus far, we have treated $\lambda$ as a fixed tuning parameter, but in practice it is preferable to have an automated, data-dependent way to choose it. Standard practice for the LASSO is to choose $\lambda$ via cross-validation (on the full data $(\bm{y},\bm{X})$), but while this choice invalidates the theoretical guarantees of the $\ell$-test, just a slight modification of it is sufficient to retain those guarantees. Let $\tilde{\bm{u}}\sim\mathrm{Unif}(\mathbb{S}^{n-d})$ be drawn independently of $\bm{u}$, and plug it into Equation (2.1) and call the resulting left-hand side $\tilde{\bm{y}}$, so that $\tilde{\bm{y}}$ is conditionally independent of $\bm{y}$ given $\bm{S}^{(j)}$. Then it is easy to see that cross-validation on $(\tilde{\bm{y}},\bm{X}_{-j})$ produces a $\lambda$, which we denote by $\hat{\lambda}_{j}$, that is exactly valid to use in the $\ell$-test, since conditioning on $\hat{\lambda}_{j}$ does not change the $\ell$-distribution [^3] Cross-validation on $(\tilde{\bm{y}},\bm{X})$ would also be valid and natural, but we prefer $(\tilde{\bm{y}},\bm{X}_{-j})$ for computational reasons; see Appendices G.1 and G.2 for details on this and other ways to choose $\lambda$ that we considered, all of which we found to be empirically dominated by $\hat{\lambda}_{j}$. And empirically, $\hat{\lambda}_{j}$ (our implementation uses 10-fold cross-validation) seems to be just as powerful as the (technically invalid) choice via cross-validation on $(\bm{y},\bm{X})$; see Appendix G.2. Although $\hat{\lambda}_{j}$ is randomized through $\tilde{\bm{u}}$ and the random 10-fold partition of the data used by cross-validation, we find this exogenous randomness barely makes a difference: in our simulations it empirically accounts for at most about 0.8% of the variability in the $\ell$-test $p$-value $p_{j}^{\hat{\lambda}_{j}}$; see Appendix G.2. Furthermore, if desired, this randomness could be arbitrarily reduced by computing many conditionally independent $\hat{\lambda}_{j}$’s and using their mean or median for the $\ell$-test.
@@ 2.5 Putting it all together: the $\ell$-test
We can now state our recommended implementation of the $\ell$-test: the $p$-value $p_{j}^{\hat{\lambda}_{j}}$ which combines the main $\ell$-test idea with the tie-breaking of Sections 2.3 and the $\lambda$ choice of Section 2.4; see Algorithm 1 for formal statement.
1Input: Data: $(\bm{y},\bm{X}_{n\times d})$, index to test: $j$
2 Output: The $\ell$-test $p$-value for $H_{j}:\beta_{j}=0$.
3 Compute $\tilde{\bm{y}}=\bm{P}_{-j}\bm{y}+\hat{\sigma}_{j}\bm{V}\tilde{\bm{u}}$, where $\tilde{\bm{u}}\sim\mathrm{Unif}\left(\mathbb{S}^{n-d}\right)$ independent of $(\bm{y},\bm{X})$.
4 Choose $\hat{\lambda}_{j}$ via 10-fold cross-validation of the LASSO on $(\tilde{\bm{y}},\bm{X}_{-j})$.
5 Return: $p_{j}^{\hat{\lambda}_{j}}$ computed by using $\lambda=\hat{\lambda}_{j}$ in Equation (2.7).
Algorithm 1 The $\ell$-test
Computing $p_{j}^{\hat{\lambda}_{j}}$ requires, aside from a cross-validated LASSO to compute $\hat{\lambda}_{j}$, one (non-cross-validated) LASSO to compute $\hat{\bm{\beta}}^{\hat{\lambda}_{j}}$ and, if $\hat{\beta}_{j}^{\hat{\lambda}_{j}}\neq 0$, then a second (non-cross-validated) LASSO to compute $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(-\hat{\beta}_{j}^{\hat{\lambda}_{j}})$. When $\hat{\beta}_{j}^{\hat{\lambda}_{j}}\neq 0$, these two LASSO’s allow us to compute the two tails for the $\ell$-test $p$-value via Theorem 1, since $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(\hat{\beta}_{j}^{\hat{\lambda}_{j}})=\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}$. And when $\hat{\beta}_{j}^{\hat{\lambda}_{j}}=0$, by Equation (2.7), the only LASSO quantity needed is $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(0)$ to compute $\hat{m}_{j}$, but since $\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}(0)=\hat{\bm{\beta}}_{-j}^{\hat{\lambda}_{j}}$ in this case, no additional LASSO run is needed beyond the first. Thus computation for the $\ell$-test requires just a very small constant number of LASSO runs.
@@
Corollary 1 (Validity of the $\ell$-test).